lf a converging lense and and diverging lense touching one another ( with light reaching coverging lense first) to act as one lense, what will be the sign (positive of negstive) for its focal lenght and its power? Explain
Answer #1

If f1 and f2 be f2 be the focal length of the convex and concave lens respectively, then according to certasian sign convention, equivalent focal length of the combination of two lenses, placed in contact with each other is given by,

1/F = 1/f1 - 1/f2 => F = f1f2/(f2 - f1)

and power of the combination, P = 1/F = 1/f1 - 1/f2

Now, the sign of the equivalent focal length and power would be positive if f1<f2 and would be negative if f1>f2 .